The sum of 2 numbers is 25. 12 less than 4 times one of the numbers is 16 more than twice the other number. Find both numbers.
Answer by Earlsdon(6103)

Let the two numbers be a and b.
a+b = 25 "The sum of two numbers is 25."
4a-12 = 2b+16 "12 less than 4 times one of the numbers (4a-12) is 16 more than twice the other number(2b+16)."
Rewrite the first equation as: a = 25-b and substitute into the second equation.
4(25-b)-12 = 2b+16 Simplify and solve for b.
100-4b-12 = 2b+16 Add 4b to both sides.
88 = 6b+16 Subtract 16 from both sides.
72 = 6b Divide both sides by 6.
12 = b and...
a = 25-b
a = 25-12
a = 13
The two numbers are 12 and 13.